8.34 A catcher catches a 145 g baseball traveling horizontally at 30.0 m/s . A. How large an impulse does the ball gi...

8.34

A catcher catches a 145 g baseball traveling horizontally at 30.0 m/s .

A. How large an impulse does the ball give to the catcher?

ANSWER: p= ___________ Kg*m/s

B. If it takes the ball 25.0 ms to stop once it is in contact with the catcher's glove, what average force did the ball exert on the catcher?

ANSWER: F= __________ N
Answer #1

A.

p=mv=(0.145kg)*(30m/s)=4.35kg.m/s

B.

F_{prom}=\left | \frac{\Delta p}{\Delta t} \right |=\left | \frac{0-4.35kg.m/s}{25*10^{-3}s} \right |=174N

Now consider that Force as impulse are vectors, but here we are working only with the magnitudes of the vector, then the magnitude of force will be positive even though \Delta p is negative.

Good luck!